OpenStudy (anonymous):

Given the equations 9x+34y=6 and 2x+12y=9, by what factor would you multiply the second equation to eliminate y and solve the system through linear combinations?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

what would you multiply \(12\) by to get \(-36\)?

OpenStudy (amorfide):

to eliminate y, you must get the same coefficient of y in both equations so since we use the second equation we want to turn 12 into 32 to do this we make an expression 12x=34 x=34/12=17/6 so we multiply by 17/6 to get the same coefficient then you can simply eliminate y